The bacteria in a certain culture double every 6.3 hours. The culture has 7,500 bacteria at the start.
How many bacteria will the culture contain after 5 hours?

Answer by jorel555(1290)

Using the formula: A = Ao*2^(t/d) where
A = amt after t time
Ao = initial amt (t=0)
t = time of action
d = doubling time of the bacteria
;
A = 7500*2^(5/6.3)
A = 7500*1.733455
A = 13000 bacteria after 5 hours!!!!!!!!!!